40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 468     Entry time: Thu May 8 01:07:24 2008
Author: rana 
Type: Summary 
Category: LSC 
Subject: Frequency Noise test: MC Trans Backscatter 
There is a wandering hump in the MC_F spectrum. It can move around on the time
scale of seconds between 40 and 200 Hz. It has an amplitude ~5-50x above the background spectrum. This seems new; I don't remember it
from a year ago. It is there in the IFO unlocked as well as the IFO locked as well as the locked + CM mode.

Tapping the AS table and/or the PSL table enclosures produces a broadband increase in the MC_F spectrum but doesn't
selectively effect the hump.

We thought it might be backscatter from the MC TRANS path and so we stuck in one of Steve's cool black glass V's into
this space. No effect. We should design a black glass V dump which we can replicate in large quantities for us and for
the sites. Something like the one on the LSC PDs, but with a 1 sq. inch opening area and a 2 inch depth.


We have also done this on the MC2 - TRANS beam before. No noise reduced there either.

The noise hump is appearing in MC_F but not in CARM_IN1 (after the CM handoff) so it seems like the MC has enough gain
to squash it. This also exonerates the MC REFL path since anything there would not be effected by the MC servo gain and
so would be visible in CARM.

My best guess is that there is something really, really scattery on the PSL table. But for now it looks like this is not a
big factor in the locking
issues.
ELOG V3.1.3-